# Standardizing Generative Face Video Compression using Supplemental Enhancement Information

### v2

@@ Abstract
This paper proposes a Generative Face Video Compression (GFVC) approach using Supplemental Enhancement Information (SEI), where a series of compact spatial and temporal representations of a face video signal (i.e., 2D/3D key-points, facial semantics and compact features) can be coded using SEI message and inserted into the coded video bitstream. At the time of writing, the proposed GFVC approach using SEI messages has been adopted into the official working draft of Versatile Supplemental Enhancement Information (VSEI) standard by the Joint Video Experts Team (JVET) of ISO/IEC JTC 1/SC 29 and ITU-T SG16, which will be standardized as a new version for “ITU-T H.274 ISO/IEC 23002-7”. To the best of the authors’ knowledge, the JVET work on the proposed SEI-based GFVC approach is the first standardization activity for generative video compression. The proposed SEI approach has not only advanced the reconstruction quality of early-day Model-Based Coding (MBC) via the state-of-the-art generative technique, but also established a new SEI definition for future GFVC applications and deployment. Experimental results illustrate that the proposed SEI-based GFVC approach can achieve remarkable rate-distortion performance compared with the latest Versatile Video Coding (VVC) standard, whilst also potentially enabling a wide variety of functionalities including user-specified animation/filtering and metaverse-related applications.
@@
Generative AI, face video coding, supplemental enhancement information, VVC, JVETI Introduction
@@
Recently, generative AI technology has greatly accelerated the progress of Generative Face Video Compression (GFVC) [1, 2, 3, 4], showing promising performance potentials and application diversities over traditional codecs [5, 6]. Different from early-day Model-Based Coding (MBC) techniques [7, 8], GFVC exploits the strong inference capabilities of deep generative models to improve the face reconstruction quality using very compact spatial and temporal representations (i.e., 2D/3D key-points, facial semantics, compact feature and others). Inspired by promising rate-distortion performance and the potential to enable more diverse applications, the Joint Video Experts Team (JVET) of ISO/IEC JTC 1/SC 29 and ITU-T SG16 has established an Ad hoc Group (AhG) for comprehensive GFVC investigations, including coordinated experimentation, software implementation, interoperability and so on. In the past year, the JVET GFVC AhG has established GFVC test conditions [9] and software tools [10, 11], and the Generative Face Video (GFV) Supplemental Enhancement Information (SEI) message has been included into the official working draft [12] of Versatile Supplemental Enhancement Information (VSEI) standard and will be finally standardized in “ITU-T H.274 ISO/IEC 23002-7”.
@@
From the perspective of JVET standardization activities, the proposed approach provides a reasonable solution to a standardize GFVC with SEI messages, bridging deep generative models with face video compression tasks using standardized bitstream format. In particular, standardizing GFVC with an SEI message allows various GFVC feature formats to be carried and combines these features seamlessly with the bitstreams from a traditional hybrid video codec without introducing normative changes to the encoder/decoder of the hybrid video codec. As a result, the proposed SEI-based GFVC standardization approach can effectively utilize advanced generative AI techniques to provide performance gains and functionalities over traditional codecs. The main contributions of this paper can be summarized as follows,
@@
The potential to standardize generative video compression with SEI messages has been explored in JVET. Taking face video as an example, a novel SEI-based standardization approach is proposed to specify compact facial representations using a unified high-level syntax structure that can be inserted into the coded video bitstream of a hybrid video codec.
@@
The proposed SEI message for generative face video can well support different facial representations, such as 2D/3D landmarks, 2D/3D keypoints, compact feature and facial semantics. Moreover, this GFV SEI message can also provide flexibility for different functionalities such as ultra-low bitrate communication, user-specified animation/filtering, and metaverse-related applications.
@@
The proposed GFV SEI message has been implemented on the top of the VVC reference software and extended with common GFVC software tools, which were both endorsed by JVET as GFVC AhG activities. As such, the entire SEI-based GFVC approach can be executed in a streamlined manner. The simulation results demonstrate the effectiveness of the proposed SEI-based GFVC approach, which can achieve promising compression performance and diverse functionalities over the VVC standard.
@@ II-A Generative Face Video Compression
In terms of ultra-low bitrate visual compression, parameterized face video compression can be tracked to early MBC techniques [7, 8] in the 1990s, which provided promising compression capability via statistical priors and image synthesis. However, limited by the model learning and generalization abilities at that time, the reconstruction quality was not sufficient for the early-day MBC techniques to be further developed and applied in the following years. Recently, inspired by the strong inference capabilities of deep generative models like Generative Adversarial Network (GAN) [13] and Diffusion Model (DM) [14], the MBC concept has been revived and further extended into a novel generative compression paradigm. Taking face video as an example, deep generative models can facilitate the characterization of a high-dimensional face signal into compact features, and use these features to perform accurate motion estimation for high-quality face video reconstruction.
In particular, keypoint-based GFVC frameworks [2] can realize ultra-low bit-rate face video communication by compressing a series of learned 2D keypoints along with their local affine transformations. A wide variety of facial representations like 2D facial landmarks [15], 3D keypoints [3], compact temporal features [4] and interactive facial semantics [16] have been proposed to support more robust and versatile GFVC algorithms. However, these algorithms still simply bridge generative animation models with compression tasks, and lack more mature consideration such as how to achieve commercial deployment and enable real applications through standardization. As for the requirements of GFVC standardization, it should focus on specifying the bitstream format for a variety of GFVC feature representations in a unified form and ensuring the interoperability/robustness of GFVC systems. Moreover, standardized GFVC format can realize ultra-low bitrate face video communication and improve the quality of experience of face video applications like video conferencing/chat. It also potentially enables a wide variety of functionalities including user-specified animation/filtering and metaverse-related applications.
@@ II-B Supplemental Enhancement Information
As additional data specified into the bitstream, SEI messages can convey extra information and support various post-processing technology. In the past decades, JVET has successfully developed many high-level syntax structures [17, 18] to enable versatile scenarios and new functionalities like viewport-adaptive 360° immersive media.
From the year of 2023, JVET has started to investigate the standardization of GFVC, covering topics from how to carry various GFVC features using standardized syntax to how to insert GFV SEI messages into the VVC coded bitstream. Specifically, a common GFV SEI message syntax [19, 20] was proposed to support the carriage of different GFVC feature formats within a VVC coded bitstream, whilst the interfaces to GFVC translator, generator, and enhancer neural networks [21, 22, 23] were also specified. Interoperability among different GFVC feature formats [24], fusion/enhancement modules for quality improvement [25, 26, 27] and model complexity reduction for practical deployment [28] were further investigated to enable GFVC applications towards better performance. GFVC software tools [10, 11] have been developed to support various GFVC algorithms, and the GFV SEI message and its enhancement extension have also been included into the official working draft [12] of VSEI standard and will be finally standardized in “ITU-T H.274 ISO/IEC 23002-7”. Herein, this paper summarizes how to enable different GFVC representations with GFV SEI message. The paper also describes a technical implementation to link GFVC software tools with GFV SEI message implementation for comprehensive performance evaluations.
@@ III SEI-based Generative Compression Approach
In this section, we provide the GFV SEI message definition and its workflow in pseudo-code format regarding how to encapsulate different GFVC feature formats with a unified GFV SEI message syntax. We also discuss how to support interoperability among different GFVC feature formats and improve reconstruction robustness via fusion/enhancement techniques. Finally, we describe the detailed encoding and decoding processes for the proposed SEI-based GFVC approach.
@@
| IFVC [16] | Facial Semantics (Mouth/Eye/Head Rotation/Translation Matrices) | Matrix |
@@ III-A GFV SEI Message Design
A face signal exhibits strong statistical regularities, which can be efficiently characterized with 2D landmarks, 2D keypoints, 3D keypoints, compact features, facial semantics or other formats. Such compact facial description strategies can lead to reduced coding bit-rate and improved coding efficiency for GFVC algorithms. However, diverse facial representations also make it difficult to standardize GFVC. Herein, we summarize a few typical GFVC algorithms and attempt to discover the characteristics of their compact representations. As shown in Table I, these compact facial representations can be classified into two categories: coordinate-related parameters and matrix-related parameters. In addition, all GFVC algorithms proposed in the literature use a hybrid codec to code the base pictures. These observations confirm that there is great potential to standardize GFVC. Intuitively, a SEI message provides a reasonable standardized way to accommodate different GFVC representations into a unified format and convey these compact representations together with VVC-coded base pictures. As such, SEI-based GFVC approach has recently been a subject active study in JVET.
Algorithm 1 provides a pseudocode representing the syntax structure of GFV SEI message according to the GFV SEI message text that has been included in the working draft of VSEI version 4 [12]. In particular, the GFV SEI message indicates facial parameters and specifies a facial parameter translator network, denoted as TranslatorNN, and a face picture generator neural network, denoted as GenerativeNN. Besides, the GFV SEI message provides specific flags/factors to determine parameter type, whether prediction is used, and codeword length. It should be emphasized that the GFV SEI message generator in Algorithm 1 is just a brief summary of the GFV SEI message included in the working draft of VSEI version 4 [12]. Readers are referred to the detailed syntax and semantics of GFV SEI message draft text in [12] for further information, including 2D/3D parameters, number of coordinates, matrix dimensions and flags for extra functionalities, as well as the usage interface for TranslatorNN and GenerativeNN that use these GFV parameters for signal reconstruction.
@@ III-B GFVC System Optimization
The proposed GFV SEI message semantics for the VSEI standard provides a specification of syntax and semantics sufficient for usage, but does not establish conformance requirements for exactly how a decoder would use the provided SEI message to produce the pictures it would display in response to receiving the SEI message. Instead, a core decoding process is applied as for ordinary picture decoding, and the GFV processing is applied as a post-process that is outside of the motion-compensated prediction loop. This approach offers decoder implementers the freedom to customize their implementations to suit their own needs or for specifications establishing such details to be developed separately in other organizations. To facilitate this approach, the draft syntax for the GFV SEI message supports the ability for an encoder to use Uniform Resource Identifiers (URIs) to identify such third-party specifications.
@@
Although the proposed GFV SEI message can encapsulate various GFVC features into a unified syntax format, the existence of various GFVC algorithms can still result in mismatched encoder and decoder, where the encoder chooses to send a type of facial representation that is different from what the generator network at the decoder expects to use. In other words, various types of GFVC features with distinct characteristics may not be mutually compatible during the face generation process. To solve the issue of interoperability among different GFVC systems, JVET has included a feature translator network [24] into the usage model for the GFV SEI message. The translator network can convert received GFVC features into ones that are specific to the decoder’s preferred model for interpretation. As such, different GFVC features can be effectively adapted to one another while retaining promising coding performance.
@@
In addition, limited by the capabilities of generative models and the internal GFVC mechanisms (e.g., unchanged textures provided by a base picture and sometimes insufficient motion representation due to compact representation), the reconstructed face video quality sometimes may not be robust. Such instability in reconstruction quality mainly includes objectionable background distortions, and/or inaccurate local motions in the mouth and/or eyes, which may damage the feasibility of practical applications. To improve reconstruction robustness, the JVET GFVC AhG has investigated a series of image processing techniques for signal fusion and enhancement, including a scalable representation and layered reconstruction scheme [35], background segmentation/inpainting/fusion operations [27, 25], and pupil position redirection [36].
III-C Encoding/Decoding Processes

### v3

@@ Abstract
This paper proposes a Generative Face Video Compression (GFVC) approach using Supplemental Enhancement Information (SEI), where a series of compact spatial and temporal representations of a face video signal (e.g., 2D/3D key-points, facial semantics and compact features) can be coded using SEI messages and inserted into the coded video bitstream. At the time of writing, the proposed GFVC approach using SEI messages has been included into a draft amendment of the Versatile Supplemental Enhancement Information (VSEI) standard by the Joint Video Experts Team (JVET) of ISO/IEC JTC 1/SC 29 and ITU-T SG21, which will be standardized as a new version of ITU-T H.274 ISO/IEC 23002-7. To the best of the authors’ knowledge, the JVET work on the proposed SEI-based GFVC approach is the first standardization activity for generative video compression. The proposed SEI approach has not only advanced the reconstruction quality of early-day Model-Based Coding (MBC) via the state-of-the-art generative technique, but also established a new SEI definition for future GFVC applications and deployment. Experimental results illustrate that the proposed SEI-based GFVC approach can achieve remarkable rate-distortion performance compared with the latest Versatile Video Coding (VVC) standard, whilst also potentially enabling a wide variety of functionalities including user-specified animation/filtering and metaverse-related applications.
@@
Generative AI, face video coding, supplemental enhancement information, VVC, VSEI, JVETI Introduction
@@
Recently, generative AI technology has greatly accelerated the progress of Generative Face Video Compression (GFVC) [1, 2, 3, 4, 5, 6], showing promising performance potentials and application diversities over traditional codecs [7, 8, 9]. Different from early-day Model-Based Coding (MBC) techniques [10, 11, 12, 13, 14], GFVC exploits the strong inference capabilities of deep generative models to improve the face reconstruction quality using very compact spatial and temporal representations (e.g., 2D/3D keypoints, facial semantics and compact features). Inspired by promising rate-distortion performance and the potential to enable more diverse applications, the Joint Video Experts Team (JVET) of ISO/IEC JTC 1/SC 29 and ITU-T SG21 (formerly SG16) has established an Ad hoc Group (AhG) for comprehensive GFVC investigations, including coordinated experimentation, software implementation, interoperability study, and so on. As illustrated in Fig. 1, over the past two and a half years, the standardization efforts of GFVC have made significant strides towards establishing a unified test environment and bitstream syntax structure for coding of face video content. In particular, the JVET GFVC AhG has established GFVC test conditions [15] and software tools [16, 17], and the Generative Face Video (GFV) and Generative Face Video Enhancement (GFVE) Supplemental Enhancement Information (SEI) messages have been included into the Versatile Supplemental Enhancement Information (VSEI) standard [18], which will be standardized as a new version of ITU-T H.274 ISO/IEC 23002-7. Also, the GFV and GFVE SEI messages have been recently added to the next version [19] of H.264/Advanced Video Coding (AVC) [7] and H.265/High Efficiency Video Coding (HEVC) [8].
@@
From the perspective of JVET standardization activities, the proposed approach provides a reasonable solution to a standardized GFVC approach using SEI messages, bridging deep generative models with face video compression tasks using a standardized bitstream format. In particular, standardizing GFVC with SEI messages allows various GFVC feature formats to be carried and combines these features seamlessly with the bitstreams from a traditional hybrid video codec (e.g., VVC, HEVC and AVC) without introducing normative changes to the encoder/decoder of the hybrid video codec. As a result, the proposed SEI-based GFVC standardization approach can effectively utilize advanced generative AI techniques to provide performance gains and functionalities over traditional codecs while retaining a form of backward compability with existing decoders. The main contributions of this paper can be summarized as follows,
@@
The potential to standardize generative video compression with SEI messages has been explored in JVET. Taking face video coding as use case, a novel SEI-based standardization approach is proposed to specify compact facial representations using a unified high-level syntax structure that can be inserted into the coded video bitstream of a hybrid video codec.
@@
The proposed SEI messages for generative face video compression can well support different facial representations, such as 2D/3D landmarks, 2D/3D keypoints, compact feature and facial semantics. Moreover, the GFV and GFVE SEI messages can also provide flexibility for different functionalities such as ultra-low bitrate communication, user-specified animation/filtering, and metaverse-related applications.
@@
The proposed GFV and GFVE SEI messages have been implemented on top of the VVC reference software (VTM) and extended with common GFVC software tools, which were both endorsed by JVET as GFVC AhG activities. As such, the entire SEI-based GFVC approach can be executed in a streamlined manner. The simulation results demonstrate the effectiveness of the proposed SEI-based GFVC approach, which can achieve promising compression performance and diverse functionalities over the VVC standard. Further, upon inclusion into the next version of the HEVC and AVC standards, the proposed SEI messages have been implemented on top of the HEVC and AVC standard reference software as well.
@@ II-A Generative Face Video Compression
In terms of ultra-low bitrate visual compression, parameterized face video compression can be tracked to early MBC techniques [10, 11, 12, 13, 14] in the 1990s, which provided promising compression capability via statistical priors and image synthesis. However, limited by the model learning and generalization abilities at that time, the reconstruction quality was not sufficient for the early-day MBC techniques to be further developed and applied in the following years. Recently, inspired by the strong inference capabilities of deep generative models like Generative Adversarial Networks (GANs) [20] and Diffusion Models (DMs) [21], the MBC concept has been revived and further extended into a novel generative compression paradigm. Taking face video as an example, deep generative models can facilitate the characterization of a high-dimensional face signal into compact features at the encoder side, and use these features to perform accurate motion estimation for high-quality face video reconstruction at the decoder side.
In particular, keypoint-based GFVC frameworks [2] can realize ultra-low bit-rate face video communication by compressing a series of learned 2D keypoints along with their local affine transformations. A wide variety of facial representations like 2D facial landmarks [22], 3D keypoints [5, 23], compact temporal features [6], interactive facial semantics [24], progressive visual tokens [4] and scalable features [25] have been proposed to support more robust and versatile GFVC algorithms. However, these algorithms still simply bridge generative animation models with compression tasks, and lack more mature consideration such as how to achieve commercial deployment and enable real applications through standardization. As for the requirements of GFVC standardization, it should focus on specifying the bitstream format for a variety of GFVC feature representations in a unified form and ensuring the interoperability/robustness of GFVC systems. Moreover, standardized GFVC format can realize ultra-low bitrate face video communication and improve the quality of experience of face video applications like video conferencing/chat. It also potentially enables a wide variety of functionalities including user-specified animation/filtering and metaverse-related applications.
@@ II-B Supplemental Enhancement Information
As additional data specified into the bitstream, SEI messages can convey extra information and support various post-processing technologies. In the past decades, JVET has successfully developed many high-level syntax features [26, 27] to enable versatile scenarios and new functionalities like viewport-adaptive 360° immersive media and film grain removal and synthesis.
From the year of 2023, JVET has started to investigate the standardization of GFVC, covering topics from how to carry various GFVC features using standardized syntax to how to insert GFV SEI messages into the VVC coded bitstream. Specifically, the first GFV SEI message, as described in [28], was introduced to enable VVC-coded pictures to serve as base pictures, supplementing them with a minor additional bit overhead to convey facial semantics. Afterwards, a common GFV SEI message syntax [29, 30] was proposed to support the carriage of different GFVC feature formats within a VVC coded bitstream, whilst the interfaces to GFVC translator, generator, and enhancer neural networks [31, 32, 33] were also specified. On top of the GFV SEI message, [34] proposed a new Generative Face Video Enhancement (GFVE) SEI message to support the scalable representation and layered reconstruction of GFVC. In addition, there were a series of syntax refinements [35, 36, 37, 30, 38] and additional operations (e.g., chroma key information [39], timing information and order [40, 41, 42], low-confidence facial motion information [43], no-display flag [44] and pupil position redirection [45]) made available to GFV and GFVE SEI messages, which have improved their syntax accuracy and enriched their functionalities.
Moreover, interoperability among different GFVC feature formats [46], fusion/enhancement modules for quality improvement [47, 48] and model complexity reduction for practical deployment [49, 50] were further investigated to enable GFVC applications towards better performance. GFVC software tools [16, 17] have been developed to support various GFVC algorithms, and the GFV/GFVE SEI messages have also been included into the working draft [18] of the VSEI standard, which will be standardized as a new version of ITU-T H.274 ISO/IEC 23002-7. Recently, the GFV and GFVE SEI messages have also been added to the next version [19] of H.264/AVC [7] and H.265/HEVC [8], which allows AVC- and HEVC-coded base pictures to be used in the GFVC framework. Herein, this paper summarizes how to enable different GFVC representations with GFV SEI message. The paper also describes a technical implementation to link GFVC software tools with GFV SEI message implementation for comprehensive performance evaluations.
@@ III SEI-based Generative Compression Approach
In this section, we provide the GFV SEI message definition and its workflow in pseudo-code format regarding how to encapsulate different GFVC feature formats with a unified GFV SEI message syntax. We also discuss how to support interoperability among different GFVC feature formats and improve reconstruction robustness via fusion/enhancement techniques. Finally, we provide implementation details of GenerativeNN/TranslatorNN, and describe the detailed encoding and decoding processes for the proposed SEI-based GFVC approach.
@@
| IFVC [24] | Facial Semantics | Matrix |
| PFVC [4] | Progressive Tokens | Matrix |
@@ III-A GFV SEI Message Design
A face signal exhibits strong statistical regularities, which can be efficiently characterized with 2D landmarks, 2D keypoints, 3D keypoints, compact features, facial semantics or other formats. Such compact facial description strategies can lead to reduced coding bit-rate and improved coding efficiency for GFVC algorithms. However, diverse facial representations also make it difficult to standardize GFVC. Herein, we summarize a few typical GFVC algorithms and attempt to discover the characteristics of their compact representations. As shown in Table I, these compact facial representations can be classified into two categories: coordinate-related parameters and matrix-related parameters. In addition, all GFVC algorithms proposed in the literature use a hybrid codec to code the base pictures. These observations confirm that there is great potential to standardize GFVC. Intuitively, a SEI message provides a reasonable standardized way to accommodate different GFVC representations into a unified format and convey these compact representations together with VVC-coded (or HEVC- or AVC-coded) base pictures. As such, SEI-based GFVC approach has recently been a subject active study in JVET.
Algorithm 1 provides a pseudocode representing the syntax structure of GFV SEI message according to the GFV SEI message text that has been included in the working draft of VSEI version 4 [18]. In particular, the GFV SEI message indicates facial parameters and specifies a facial parameter translator network, denoted as TranslatorNN, and a face picture generator neural network, denoted as GenerativeNN. Besides, the GFV SEI message provides specific flags/factors to determine parameter type, whether prediction is used, and codeword length. It should be emphasized that the GFV SEI message in Algorithm 1 is just a brief summary of the GFV SEI message included in the working draft of VSEI version 4 [18]. Readers are referred to the detailed syntax and semantics of GFV SEI message draft text in [18] for further information, including 2D/3D parameters, number of coordinates, matrix dimensions and flags for extra functionalities, as well as the usage interface for TranslatorNN and GenerativeNN that use these GFV parameters for signal reconstruction.
@@ III-B GFVC System Optimization
The proposed GFV SEI message semantics for the VSEI standard provides a specification of syntax and semantics sufficient for usage, but does not establish conformance requirements for exactly how a decoder would use the provided SEI message to produce the pictures it would display in response to receiving the SEI message. Instead, a core decoding process is applied as for ordinary picture decoding to generate the base pictures, and the GFV processing is applied as a post-process that is outside of the motion-compensated prediction loop. This approach offers decoder implementers the freedom to customize their implementations to suit their own needs or for specifications establishing such details to be developed separately in other organizations. To facilitate this approach, the draft syntax for the GFV SEI message supports the ability for an encoder to use Uniform Resource Identifiers (URIs) to identify such third-party specifications.
@@
Although the proposed GFV SEI message can encapsulate various GFVC features into a unified syntax format, the existence of various GFVC algorithms can still result in mismatched encoder and decoder, where the encoder chooses to send a type of facial representation that is different from what the generator network at the decoder expects to use. In other words, various types of GFVC features with distinct characteristics may not be mutually compatible during the face generation process. To solve the issue of interoperability among different GFVC systems, JVET has included a feature translator network [46, 57] into the usage model for the GFV SEI message. The translator network can convert received GFVC features into ones that are specific to the decoder’s preferred model for interpretation. As such, different GFVC features can be effectively adapted to one another while retaining promising coding performance.
@@
In addition, limited by the capabilities of generative models and the internal GFVC mechanisms (e.g. 1) if the base picture is not updated, the texture reference for subsequent picture generation will be unchanged; and 2) compact representations maintain a fixed dimensionality regardless of scene variations, sometimes limiting their ability to capture complex motion patterns in dynamic environments.), the reconstructed face video quality sometimes may not be stable. Such instability in reconstruction quality mainly includes objectionable background distortions, and/or inaccurate local motions in the mouth and/or eyes, which may damage the feasibility of practical applications. To improve reconstruction robustness, the JVET GFVC AhG has investigated a series of image processing techniques for signal fusion and enhancement, including a scalable representation and layered reconstruction scheme [47], background segmentation/inpainting/fusion operations [48, 39], pupil position redirection [45], colour calibration [58] and QP-adaptive feature enhancement [59].
III-C Implementations of GenerativeNN and TranslatorNN
As shown in Fig. 3.(a), GenerativeNN specifies a neural network that may be used to generate output pictures using the fixed format of facial parameters and previously decoded output pictures. It should be emphasized that GenerativeNN does not employ a fixed network architecture. Instead, it refers to any generative model capable of synthesizing facial images with compact facial parameters. Taking CFTE/FOMM/DAC/FV2V as examples, GenerativeNN commonly consists of two components: a dense motion estimator and a generator. The dense motion estimator converts decoded facial parameters into dense motion fields, while the generator further combines the previously decoded output pictures with these transformed dense motion fields to produce the final synthesized face results.
As shown in Fig. 3.(b), TranslatorNN specifies a neural network that may be used to convert various formats of facial parameters signaled in the SEI message into a fixed format of parameters. As such, TranslatorNN could achieve the seamless translation between different facial representations and avoid the mismatch problem bewteen GFVC encoder and decoder. TranslatorNN is designed in a portable and light-weighted manner, which consists of a trans-encoder and a trans-decoder. In particular, each trans-en/decoder consists of 4 full-connected (FC) layers with ReLU activation in between, and the dimension of FC layers and latent embedding is set to 256.
III-D Encoding/Decoding Processes